# A Simple and Efficient Joint Measurement Strategy for Estimating Fermionic Observables and Hamiltonians

### v2

@@ Abstract
We propose a simple scheme to estimate fermionic observables and Hamiltonians relevant in quantum chemistry and correlated fermionic systems. Our approach is based on implementing a measurement that jointly measures noisy versions of any product of two or four Majorana operators in an $N$ mode fermionic system. To realize our measurement we use: (i) a randomization over a set of unitaries that realize products of Majorana fermion operators; (ii) a unitary, sampled at random from a constant-size set of suitably chosen fermionic Gaussian unitaries; (iii) a measurement of fermionic occupation numbers; (iv) suitable post-processing. Our scheme can estimate expectation values of all quadratic and quartic Majorana monomials to $\epsilon$ precision using $\mathcal{O}(N\log(N)/\epsilon^{2})$ and $\mathcal{O}(N^{2}\log(N)/\epsilon^{2})$ measurement rounds respectively, matching the performance offered by fermionic shadow tomography [1, 2]. In certain settings, such as a rectangular lattice of qubits which encode an $N$ mode fermionic system via the Jordan-Wigner transformation, our scheme can be implemented in circuit depth $\mathcal{O}(N^{1/2})$ with $\mathcal{O}(N^{3/2})$ two-qubit gates, offering an improvement over fermionic and matchgate classical shadows that require depth $\mathcal{O}(N)$ and $\mathcal{O}(N^{2})$ two-qubit gates. By benchmarking our method on exemplary molecular Hamiltonians and observing performances comparable to fermionic classical shadows, we demonstrate a novel, competitive alternative to existing strategies.
@@ II.1 Setting and notation
with $\eta_{A}\in(0,1]$ the visibility and $e_{A}\in\{\pm 1\}$ the measurement outcome (see Fig. 1 for an illustration of the scheme).
@@ II.2 A General Joint Measurement Strategy
Figure 1: Circuit implementation of the joint measurement scheme described in (i)–(iv) in Sec. II.2 for an $N$ mode fermionic state. In step (i) the state evolves under the action of a Majorana monomial $\gamma_{X}^{\dagger}$, sampled uniformly at random from the set of $2^{2N}$ Majorana monomials. In step (ii) a fermionic Gaussian unitary $U^{\dagger}$ is applied. In step (iii) the $N$ projective measurements of the disjoint Majorana pairs are performed, with outcomes $q_{1},\ldots,q_{N}$. Finally, for a given $A\subset[2N]$, one outputs $e_{A}$ according to step (iv).
@@ II.3 Gaussian unitaries for measurements of Majorana pairs and quadruples
Figure 2: The $202N=$ 20 Majorana operators are partitioned into 5 subsets (rows) $Y^{(r)}_{\alpha}$ (represented by light blue rounded rectangles) of cardinality 4, for each measurement setting $r=1,2$. In the first measurement setting (a), the unitary $U_{\text{sup}}$ acts via a lower flat orthogonal matrix on each row $Y_{\alpha}^{(1)}$, transforming every $\gamma_{k}$, $k\in Y_{\alpha}^{(1)}$, into a balanced superposition $\tilde{\gamma}_{k}=\sum_{l\in Y_{\alpha}^{(1)}}b^{(\alpha)}_{lk}\gamma_{l}$, with $b^{(\alpha)}_{lk}=\Omega(N^{-1/4})$. The pairings $(\pi(2j-1),\pi(2j))$ (marked by solid black lines) are constructed by permuting the standard pairings $\{(2j-1,2j)\}_{j=1}^{N}$ (represented by dark blue rounded rectangles) with a permutation $\pi$ that vertically shifts (cyclically) the second and fourth columns by 1 and 2 respectively. These pairings ensure the sample from a noisy variant of $\gamma_{ij}$ can be obtained if $i$ and $j$ are in different rows. In the second setting (b), the scheme is modified by reshuffling the operators by a permutation $\sigma^{(2)}$, that shifts columns 2,3,4 (of setting 1) vertically downwards by 1,2,3 rows, respectively. All operators in the same row of setting 1 now appear in different rows of setting 2, ensuring joint measurability of all pairs. The standard pairings represented by dark blue rectangles are consistent with the JW encoding in a 2D grid (see Fig. 4), where each such rectangle corresponds to one qubit.
@@
This modification realizes a joint measurement of all pairs and quadruples $A$ with visibilities $\eta_{ij}\geq(1/K_{2})\tilde{\eta}_{ij}$ and $\eta_{ijkl}\geq(1/K_{4})\tilde{\eta}_{ijkl}$, where $\tilde{\eta}_{A}$ is the visibility attainable from the unitary whose partition is consistent with $A$, and $K_{2},K_{4}$ are the minimal numbers of measurement settings that ensure all pairs and quadruples are compatible with at least one measurement round, respectively. In Fig. 2 we give an example of two measurement settings ($K_{2}=2$) that ensure all pairs are measured for a system of $2N=L(L+1)$ Majorana modes arranged in a rectangular layout (with $L$–even). It can also be shown that when $L+1$ is a prime number, all Majorana quadruples can be measured using $K^{\text{prime}}_{4}=7$ settings via a method that is compatible with the original layout and can be obtained by permuting Majorana modes in columns of the Majorana fermion arrangement from Fig. 2 (part a). See Appendix A for the proof. For general $L$, a probabilistic reasoning given in [36] ensures that $K_{4}=9$ random partitions of the $2N=L(L+1)$ modes into subsets of size $L$ (not necessarily compatible with the 2D layout) suffice to measure all quadruples. Furthermore, we can extend this to arbitrary $N$ mode systems by relaxing the constraint that the subsets are equally sized.
@@ II.4 Strategy for physical fermionic Hamiltonians
Figure 3: Schematic representation of the construction of the 4 rounds of our measurement strategy for physical fermionic Hamiltonians from Sec. II.4. The $2N$ Majorana modes are arranged on an $2L\times L$ grid and partitioned into $L$ even ($E_{\alpha}$) and $L$ odd ($O_{\alpha}$) subsets (clusters) of cardinality $L$. The unitary $U_{\text{sup}}$ acts via a lower-flat orthogonal matrix on each cluster. In round 1, the pairings $(\pi(2j-1),\pi(2j))$ which define $U_{\text{pair}}$ (and marked by solid black lines) are constructed by a permutation $\pi$ that shifts the second and third even column cyclically by 1 and 2 respectively (in a manner analogous to Fig. 2). Importantly, the pairings (black lines) ensure each even cluster is coupled to each odd cluster. In rounds $r=2,3,4$, the different partitions of the modes into even and odd clusters are implemented by $U^{(r)}_{\text{resh}}$ via the permutations $\sigma^{(r)}$. The permutation $\sigma^{(2)}$ is realized by a vertical cyclical shift of the second and third odd columns of the first measurement setting by 1 and 2, respectively (bottom left). Similarly, $\sigma^{(3)}$ corresponds to a vertical cyclical shift of the second and third even columns of the first measurement (bottom middle). Finally, $\sigma^{(4)}$ is realized by performing both the vertical shifts realizing $\sigma^{(2)}$ and $\sigma^{(3)}$ (bottom right).
@@
In Appendix B we show that these four unitaries $U^{(r)}$ are sufficient to estimate all quadratic and quartic terms in $H$ with visibilities from Prop. 2. In particular, for $A\in\mathcal{X}_{2}\cup\mathcal{X}_{4}$, there exists at least one measurement round $r$ in which $A\in\mathcal{M}^{(r)}$ and therefore $\eta^{(r)}_{A}=\Theta(N^{-|A|/4})$. Note that because the FLO unitaries appearing in points (a)-(c) are either horizontal or vertical (see (12) for definition), they can be implemented with the same gate depth and gate count as those discussed in Sec. II.3.
@@
and $e^{(r)}_{A}\in\{\pm 1\}$ is the outcome obtained from the post-processing in step (iv) of the general protocol with unitary $U^{(r)}$ (cf. Sec. II.2). We choose coefficients $\alpha^{(r)}\in\mathbb{R}$ satisfying $\sum_{r:A\in\mathcal{M}^{(r)}}\alpha_{A}^{(r)}=1$, which ensures $\mathbb{E}\left[\sum_{r}\alpha_{A}^{(r)}\hat{\gamma}_{A}^{(r)}\right]=\mathrm{tr}(\gamma_{A}\rho)$ and therefore $\hat{H}$ is an unbiased estimator of $\mathrm{tr}(H\,\rho)$. For details of the post-processing see Appendix C. For details of the coefficient choice see Appendix E. To improve the efficiency of our estimation strategy we can apply the median-of-means method (in analogy with classical shadows [24]). This provides a simple classical post-processing in order to reduce the effect of estimation errors, and the sample complexity is bounded by a multiple of the variance $\mathrm{Var}[\hat{H}]$ of the estimator, which for our scheme is given explicitly in Appendix D.
@@ II.5 Numerical benchmarks
We benchmark our strategy by calculating the variance of our estimator on the ground state of molecular Hamiltonians that are commonly used in the literature (see e.g. [1, 2, 25, 35, 28, 30, 44]). Table 1 presents a comparison between the variances resulting from our strategy and other measurement approaches, including fermionic classical shadows [1].
@@
Table 1: Comparison of variances for estimating energies of standard benchmark molecules. The variances are calculated for the ground states of five molecular Hamiltonians, with units Ha2. The first four strategies (taken from a summary in [1]) include: classical shadows with Pauli measurements (CS-Pauli [24]); locally biased classical shadows (LBCS [25]); basis-rotation grouping (BRG [35]); and classical shadows with fermionic Gaussian unitaries (CS-FGU [1]). Where applicable ([24, 25, 35]), the Jordan-Wigner encoding was used. The last column summarises our results, optimized over the coefficients $\alpha_{A}^{(r)}$.
For the calculations in Table 1, the coefficients $\alpha_{A}^{(r)}$ of the estimator (14) are chosen as optimal (i.e., minimizing the variance) for the ground state of a given Hamiltonian. However state-independent coefficients, uniform over rounds, can also be chosen. For details of the optimization and coefficient choice see Appendix E. Note that our estimator requires four distinct measurement rounds, and a fair comparison with strategies using a single round requires multiplying our variance by four. For further technical details see Appendix F.
By taking advantage of the fact that FLO circuits can be efficiently classically simulated [38, 45], we also implemented a simulation of our strategy on Gaussian states, whose energies can be easily calculated analytically (see e.g. [46]). We used the package FLOYao.jl [47] to classically simulate individual runs of our protocol, verifying (up to statistical precision) that the strategy correctly estimates the above molecular Hamiltonians on random Gaussian states. The empirical variance of this estimation also confirms (within statistical precision) the accuracy of the variance formula derived in Appendix D.
@@ III Discussion
There are several potential directions in which this work can be developed further. One interesting question to consider is how the choice of unitaries in the measurement process, discussed in Sec. II.3, can be optimized for a particular Hamiltonian, in order to estimate the relevant products with higher visibility. The choice of unitary can significantly affect the variance, as demonstrated in Table 2 of Appendix E. Optimization methods have already led to substantial reductions in the sample complexity of classical shadow protocols, especially in the locally biased regime (see Table 1). Moreover, in this work we only focus on degree two and four Majorana operators. A natural extension would be to consider joint measurement strategies for higher degree Majorana observables, as well as to investigate whether other layouts, fermion-to-qubit encodings, or more general orthogonal transformations, provide performance benefits. Furthermore, the effect of noisy physical implementations on the scheme could be studied, similarly to [30]. Finally, our strategy can be directly applied on native fermionic computing or simulation platforms, e.g. programmable neutral atom arrays [49], without the need to map to qubits, and is naturally compatible with devices using a 2D layout. An interesting research direction would be to experimentally apply it on such devices.

### v3

# SupplementaryInformation

@@ Abstract
We propose a simple scheme to estimate fermionic observables and Hamiltonians relevant in quantum chemistry and correlated fermionic systems. Our approach is based on implementing a measurement that jointly measures noisy versions of any product of two or four Majorana operators in an $N$ mode fermionic system. To realize our measurement we use: (i) a randomization over a set of unitaries that realize products of Majorana fermion operators; (ii) a unitary, sampled at random from a constant-size set of suitably chosen fermionic Gaussian unitaries; (iii) a measurement of fermionic occupation numbers; (iv) suitable post-processing. Our scheme can estimate expectation values of all quadratic and quartic Majorana monomials to $\epsilon$ precision using $\mathcal{O}(N\log(N)/\epsilon^{2})$ and $\mathcal{O}(N^{2}\log(N)/\epsilon^{2})$ measurement rounds respectively, matching the performance offered by fermionic classical shadows [1, 2]. In certain settings, such as a rectangular lattice of qubits which encode an $N$ mode fermionic system via the Jordan-Wigner transformation, our scheme can be implemented in circuit depth $\mathcal{O}(N^{1/2})$ with $\mathcal{O}(N^{3/2})$ two-qubit gates, offering an improvement over fermionic and matchgate classical shadows that require depth $\mathcal{O}(N)$ and $\mathcal{O}(N^{2})$ two-qubit gates. By benchmarking our method on exemplary molecular Hamiltonians and observing performances comparable to fermionic classical shadows, we demonstrate a novel, competitive alternative to existing strategies.
@@ I Introduction
We analyze the 2D rectangular lattice as both a simple theoretical model and with practical applications in mind. Many present-day superconducting quantum computing architectures are based on 2D lattices that can be easily mapped to rectangular ones. For example, IBM’s hexagon-heavy lattices can simulate a rectangular lattice with constant overhead in gate depth and count [36], and a rectangular lattice can be embedded into the topology of the Google Sycamore processor.
@@ II.1 Setting and notation
with $\eta_{A}\in(0,1]$ the visibility and $e_{A}\in\{\pm 1\}$ the measurement outcome (see Fig. 1 for an illustration of the scheme). In our work, unless otherwise stated, the term noisy refers to unsharp (non-projective) measurements. The “noise” arises from a classical post-processing of the joint measurement rather than from an imperfect measurement device.
@@
Figure 1: Circuit implementation of the joint measurement scheme for an $N$ mode fermionic state.
The scheme is described in (i)–(iv) in Sec. II.2. In step (i) the state $\rho$ evolves under the action of a Majorana monomial $\gamma_{X}^{\dagger}$, sampled uniformly at random from the set of $2^{2N}$ Majorana monomials. In step (ii) a fermionic Gaussian unitary $U^{\dagger}$ is applied. In step (iii) the $N$ projective measurements of the disjoint Majorana pairs are performed, with outcomes $q_{1},\ldots,q_{N}$. Finally, for a given $A\subset[2N]$, one outputs $e_{A}$ according to step (iv).
@@ II.3 Gaussian unitaries for measurements of Majorana pairs and quadruples
Figure 2: Example of two measurement settings for measuring all Majorana pairs for $N=10$.
The $202N=$ 20 Majorana operators are partitioned into 5 subsets (rows) $Y^{(r)}_{\alpha}$ (represented by light blue rounded rectangles) of cardinality 4, for each measurement setting $r=1,2$. In the first measurement setting (a), the unitary $U_{\text{sup}}$ acts via a lower flat orthogonal matrix on each row $Y_{\alpha}^{(1)}$, transforming every $\gamma_{k}$, $k\in Y_{\alpha}^{(1)}$, into a balanced superposition $\tilde{\gamma}_{k}=\sum_{l\in Y_{\alpha}^{(1)}}b^{(\alpha)}_{lk}\gamma_{l}$, with $b^{(\alpha)}_{lk}=\Omega(N^{-1/4})$. The pairings $(\pi(2j-1),\pi(2j))$ (marked by solid black lines) are constructed by permuting the standard pairings $\{(2j-1,2j)\}_{j=1}^{N}$ (represented by dark blue rounded rectangles) with a permutation $\pi$ that vertically shifts (cyclically) the second and fourth columns by 1 and 2 respectively. These pairings ensure the sample from a noisy variant of $\gamma_{ij}$ can be obtained if $i$ and $j$ are in different rows. In the second setting (b), the scheme is modified by reshuffling the operators by a permutation $\sigma^{(2)}$, that shifts columns 2,3,4 (of setting 1) vertically downwards by 1,2,3 rows, respectively. All operators in the same row of setting 1 now appear in different rows of setting 2, ensuring joint measurability of all pairs. The standard pairings represented by dark blue rectangles are consistent with the JW encoding in a 2D grid (see Fig. 4), where each such rectangle corresponds to one qubit.
@@
This modification realizes a joint measurement of all pairs and quadruples $A$ with visibilities $\eta_{ij}\geq(1/K_{2})\tilde{\eta}_{ij}$ and $\eta_{ijkl}\geq(1/K_{4})\tilde{\eta}_{ijkl}$, where $\tilde{\eta}_{A}$ is the visibility attainable from the unitary whose partition is consistent with $A$, and $K_{2},K_{4}$ are the minimal numbers of measurement settings that ensure all pairs and quadruples are compatible with at least one measurement round, respectively. In Fig. 2 we give an example of two measurement settings ($K_{2}=2$) that ensure all pairs are measured for a system of $2N=L(L+1)$ Majorana modes arranged in a rectangular layout (with $L$–even). It can also be shown that when $L+1$ is a prime number, all Majorana quadruples can be measured using $K^{\text{prime}}_{4}=7$ settings via a method that is compatible with the original layout and can be obtained by permuting Majorana modes in columns of the Majorana fermion arrangement from Fig. 2 (part a). See Supplementary Information A for the proof. For general $L$, a probabilistic reasoning given in [37] ensures that $K_{4}=9$ random partitions of the $2N=L(L+1)$ modes into subsets of size $L$ (not necessarily compatible with the 2D layout) suffice to measure all quadruples. Furthermore, we can extend this to arbitrary $N$ mode systems by relaxing the constraint that the subsets are equally sized.
@@
Figure 3: Schematic representation of the construction of the 4 rounds of our measurement strategy for physical fermionic Hamiltonians.
The Hamiltonians are described in Sec. II.4. The $2N$ Majorana modes are arranged on an $2L\times L$ grid and partitioned into $L$ even ($E_{\alpha}$) and $L$ odd ($O_{\alpha}$) subsets (clusters) of cardinality $L$. The unitary $U_{\text{sup}}$ acts via a lower-flat orthogonal matrix on each cluster. In round 1, the pairings $(\pi(2j-1),\pi(2j))$ which define $U_{\text{pair}}$ (and marked by solid black lines) are constructed by a permutation $\pi$ that shifts the second and third even column cyclically by 1 and 2 respectively (in a manner analogous to Fig. 2). Importantly, the pairings (black lines) ensure each even cluster is coupled to each odd cluster. In rounds $r=2,3,4$, the different partitions of the modes into even and odd clusters are implemented by $U^{(r)}_{\text{resh}}$ via the permutations $\sigma^{(r)}$. The permutation $\sigma^{(2)}$ is realized by a vertical cyclical shift of the second and third odd columns of the first measurement setting by 1 and 2, respectively (bottom left). Similarly, $\sigma^{(3)}$ corresponds to a vertical cyclical shift of the second and third even columns of the first measurement (bottom middle). Finally, $\sigma^{(4)}$ is realized by performing both the vertical shifts realizing $\sigma^{(2)}$ and $\sigma^{(3)}$ (bottom right).
@@ II.4 Strategy for physical fermionic Hamiltonians
In Supplementary Information B we show that these four unitaries $U^{(r)}$ are sufficient to estimate all quadratic and quartic terms in $H$ with visibilities from Prop. 2. In particular, for $A\in\mathcal{X}_{2}\cup\mathcal{X}_{4}$, there exists at least one measurement round $r$ in which $A\in\mathcal{M}^{(r)}$ and therefore $\eta^{(r)}_{A}=\Theta(N^{-|A|/4})$. Note that because the FLO unitaries appearing in points (a)-(c) are either horizontal or vertical (see (12) for the definition), they can be implemented with the same gate depth and gate count as those discussed in Sec. II.3.
@@
and $e^{(r)}_{A}\in\{\pm 1\}$ is the outcome obtained from the post-processing in step (iv) of the general protocol with unitary $U^{(r)}$ (cf. Sec. II.2). We choose coefficients $\alpha^{(r)}\in\mathbb{R}$ satisfying $\sum_{r:A\in\mathcal{M}^{(r)}}\alpha_{A}^{(r)}=1$, which ensures $\mathbb{E}\left[\sum_{r}\alpha_{A}^{(r)}\hat{\gamma}_{A}^{(r)}\right]=\mathrm{tr}(\gamma_{A}\rho)$ and therefore $\hat{H}$ is an unbiased estimator of $\mathrm{tr}(H\,\rho)$. For details of the post-processing see Supplementary Information C. For details of the coefficient choice see Supplementary Information E. To improve the efficiency of our estimation strategy we can apply the median-of-means method (in analogy with classical shadows [24]). This provides a simple classical post-processing in order to reduce the effect of estimation errors, and the sample complexity is bounded by a multiple of the variance $\mathrm{Var}[\hat{H}]$ of the estimator, which for our scheme is given explicitly in Supplementary Information D.
@@ II.5 Numerical benchmarks
Table 1: Comparison of variances of a single measurement round for estimating energies of standard benchmark molecules.
The variances are calculated for the ground states of five molecular Hamiltonians, with units Ha2. The first four strategies (taken from a summary in [1]) include: classical shadows with Pauli measurements (CS-Pauli [24]); locally biased classical shadows (LBCS [25]); basis-rotation grouping (BRG [35]); and classical shadows with fermionic Gaussian unitaries (CS-FGU [1]). Where applicable ([24, 25, 35]), the Jordan-Wigner encoding was used. The last column summarizes our results, optimized over the coefficients $\alpha_{A}^{(r)}$. To ensure a fair comparison of the sample complexities, the variances in the last column should be multiplied by a factor of four to account for the need of four independent measurement rounds to estimate the energy.
We benchmark our strategy by calculating the variance of our estimator (using formula (25)) on the ground state of molecular Hamiltonians that are commonly used in the literature (see e.g. [1, 2, 25, 35, 28, 30, 45]) and allow for a comparison with a wide range of strategies. Table 1 presents a comparison between the variances resulting from our strategy and other measurement approaches, including fermionic classical shadows [1]. It is standard in benchmarks [25, 1] to examine the variance as an indicator of the effectiveness of a strategy.
For the calculations in Table 1, the coefficients $\alpha_{A}^{(r)}$ of the estimator (14) are chosen as optimal (i.e., minimizing the variance) for the ground state of a given Hamiltonian. The optimized coefficients are only applicable if prior information about the state and Hamiltonian is known. However, state-independent coefficients—uniform over rounds—can also be chosen, and they perform comparatively for larger molecules. For details of the optimization and coefficient choice, see Supplementary Information E. Note that our estimator requires four distinct measurement rounds. Using standard concentration arguments, the sample complexity is bounded by a multiple of the variance [24], by e.g. Hoeffding’s inequality. Thus a fair comparison of sample complexities against strategies using single round estimators requires multiplying our variance by four. For further technical details on the numerical benchmarks see Supplementary Information F.
To analyze the results of Table 1, we first note that BRG and LBCS are strategies tailored to specific Hamiltonians. It can be expected that for more complex Hamiltonians, and larger system sizes, their advantage would diminish, due to the large number of terms in the Hamiltonian’s decomposition, making the bias towards certain observables in the decomposition less beneficial. Moreover, the optimized strategies require numerical optimizations, prior knowledge of the Hamiltonian, and a classical approximation of the underlying quantum state. For these strategies, experiments are tailored towards a specific Hamiltonian, and therefore estimating other observables with the same data is not expected to perform well. This is not the case for non-optimized strategies, including our scheme. The optimization we propose is optional and performed only in post-processing, hence it can be performed independently on each new observable of interest.
By taking advantage of the fact that FLO circuits can be efficiently classically simulated [39, 46], we also implemented a simulation of our strategy on Gaussian states, whose energies can be easily calculated analytically (see e.g. [47]). For details, see Supplementary Information G. We used the package FLOYao.jl [48] to classically simulate individual runs of our protocol, verifying (up to statistical precision) that the strategy correctly estimates the above molecular Hamiltonians on random Gaussian states. The empirical variance of this estimation also confirms (within statistical precision) the accuracy of the variance formula derived in Supplementary Information D.
@@ III Discussion
We expect our method to be suitable for NISQ devices. In the 2D setting, the gate scaling advantage is expected to become apparent for system sizes of the order of hundreds of qubits due to the overhead in gate depth and count incurred by implementing the vertical FLO transformations (see discussion in Section IV.2). Moreover, for these system sizes, ease of error mitigation for our method persists, no classical optimization is necessary, and optimized methods are expected to yield less advantage for more complex Hamiltonians. Finally, in current superconducting architectures, implementing arbitrary randomized circuits quickly (especially those required by fermionic classical shadows) is a non-trivial challenge. In contrast, randomized single-layer Paulis can be effectively implemented due to the prevalence of randomized benchmarking and mitigation methods.
There are several potential directions in which this work can be developed further. One interesting question to consider is how the choice of unitaries in the measurement process, discussed in Sec. II.3, can be optimized for a particular Hamiltonian, in order to estimate the relevant products with higher visibility. The choice of unitary can significantly affect the variance, as demonstrated in Table 2 of Supplementary Information E. Optimization methods have already led to substantial reductions in the sample complexity of classical shadow protocols, especially in the locally biased regime (see Table 1). Moreover, in this work we only focus on degree two and four Majorana operators. A natural extension would be to consider joint measurement strategies for higher degree Majorana observables, as well as to investigate whether other layouts, fermion-to-qubit encodings, or more general orthogonal transformations, provide performance benefits. Furthermore, the effect of noisy physical implementations on the scheme could be studied, similarly to [30]. Finally, our strategy can be directly applied on native fermionic computing or simulation platforms, e.g. programmable neutral atom arrays [49], without the need to map to qubits, and is naturally compatible with devices using a 2D layout. An interesting research direction would be to experimentally apply it on such devices.